OpenStudy (anonymous):

Any help with this? Using complete sentences, explain how to convert f(x)=3(x-11)2+1 into standard form.

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

square \(x-11\) first by the distributive property \[(x-11)^2=(x-11)(x-11)=...\]

OpenStudy (anonymous):

multiply the result by \(3\)

OpenStudy (anonymous):

then add 1

OpenStudy (anonymous):

is that all?

OpenStudy (anonymous):

@satellite73

OpenStudy (usukidoll):

yeah just expand (x-11)^2 first . then multiply by 3 in that portion and add 1 your standard form should look like ax^2+bx+c which is the quadratic equation

OpenStudy (usukidoll):

3(x-11)(x-11)+1 expand the (x-11)(x-11) using FOIL distribute the 3 after using FOIL for (x-11)(x-11) add the 1